# Stability of Non-Hyperbolic Sets

### v1

@@ 5.1 Continuation of secondary tangencies: the implicit function theorem
$D\Psi({0,0,0})=\left(\begin{matrix}0&\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}&\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\\ \left(\frac{\mu^{3}}{\lambda^{2-2\theta}}\right)^{n}\Psi_{2,1}&\Psi_{2,2}&\Psi_{2,3}\end{matrix}\right).$ (5.1)
In particular, by (4.2), (2.9), (4.2), (4.17) (in the estimation of the order term)
$\Psi_{1,2}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial t}\right]$ (5.2)
$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+DC^{2}n\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)+O\left(n\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\right)-\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\frac{\partial W_{n-n_{0}}}{\partial t}$
$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}\left\{{DC^{2}}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]-\frac{1}{n\left(\lambda^{\theta}\mu\right)^{2n}}\frac{\partial W_{n-n_{0}}}{\partial t}\right\}$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)=DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+n\left(\lambda^{\theta}\mu\right)^{n}v+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)$
$=$ $DC\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(n\left(\lambda^{\theta}\mu\right)^{n}\right),$
and by (4.2), (2.9), (4.2), Lemma 2 and (4.17) (in the estimation of the order term)
$\Psi_{1,3}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}\right]$ (5.3)
$=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left\{DC\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
$+$ $\left.DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right$.
$+$ $\left.O\left(\mu^{n}\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)+O\left(\frac{n-n_{0}}{\mu^{n-n_{0}}}\right)\right\}$
$=$ $DC\left[1+O\left({\left({\lambda}^{1+\theta-\alpha}{\mu}\right)^{n}}\right)\right].$
@@
Moreover $\Psi_{1,2},\Psi_{1,3}>0$ and $\Psi_{2,1}\neq 0$, see Lemma 12, (4.2) and (2.10). Because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$. Moreover the tangent space of $\Psi$ in zero satisfies $T_{(0,0,0)}\Psi^{-1}(0)=\text{Ker}D\Psi$. Hence
@@
$\frac{db}{dt}=-\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}=-\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}-{n\lambda^{\theta n}}v+O\left(\lambda^{\theta n}\right).$ (5.4)
In particular $db/dt$ is a negative number and $b$ is the graph of a smooth function. Furthermore, by the previous estimate on the slope of $b$, by Proposition 3 and by (2.11), the curve $b$ have a angle satisfying the following,
$\frac{db}{dt}=\frac{da_{n}}{dt}-{n\lambda^{\theta n}}v+O\left(\lambda^{\theta n}\right),$ (5.5)
with $v$ a positive constant.
Let $(t_{0},a_{0})$ be any parameter in $\mathcal{B}_{n}$ corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$. The proof is exactly the same as the previous case. In particular, by (4.2), (2.9), (4.2), (4.17) (in the estimation of the order term)
$\Psi_{1,2}$ $\geq$ $D(C-E))\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(n\left(\lambda^{\theta}\mu\right)^{n}\right),$
and by (4.2), (2.9), (4.2), Lemma 2 and (4.17) (in the estimation of the order term)
$\Psi_{1,3}$ $\geq$ $D(C-E)\left[1+O\left({\left({\lambda}^{1+\theta-\alpha}{\mu}\right)^{n}}\right)\right].$
Moreover, by (4.43), (2.9), (4.2),
$\Psi_{1,2}$ $\leq$ $DT\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right),$
and by (4.2), (2.9), (4.2), Lemma 2
$\Psi_{1,3}$ $\leq$ $DT\left[1+O\left(\frac{1}{\mu^{n-n_{0}/2}}\right)\right].$
Since, by Proposition 2 and (F6), $D,C-E,T,{\partial\mu}/{\partial t}>0$, then $\Psi_{1,2}>0$, $\Psi_{1,3}>0$ and $\Psi_{2,1}>0$, see Lemma 12. As before, because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$. Moreover
$\frac{(C-E)}{T}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(n\lambda^{\theta n}\right)\leq\frac{db}{dt}\leq-\frac{T}{(C-E)}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right).$ (5.6)
The proof is completed.
5.2 Limitations to the existence of secondary tangencies
We prove now that for any $n_{0}$ by moving in the direction of the $a$ parameter one can find only one tangency of type $n_{0}$ as well as when one move on the $a_{n}$ curve.
Lemma 14.
@@ 5.2 Limitations to the existence of secondary tangencies
$\frac{\partial z_{3,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}>0$
and
$\frac{\partial z_{3,y}}{\partial t}+\frac{\partial z_{3,y}}{\partial a}\frac{da_{n}}{dt}>0.$
@@
Using the estimates in Proposition 2, one can notice that the first coefficient of $\partial z_{3,y}/{\partial a}$ is strictly larger than one, while the first coefficient of $\partial W_{n-n_{0}}/{\partial a}$ is strictly smaller than one, see Lemma 2. The first inequality follows.
For the second one, if $(t,a)\in\mathcal{B}_{n}$, by Proposition 2 and (2.11),
$\frac{\partial z_{3,y}}{\partial t}+\frac{\partial z_{3,y}}{\partial a}\frac{da_{n}}{dt}=DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$
The statement is proved using Proposition 2 and (2.18). ∎
@@
Lemma 15.
There are no curve secondary tangencies of type $n_{0}$ along the line $a_{n}-E{\lambda^{\theta n}}$ where $E=C-\frac{\epsilon}{\mu^{n_{0}/2}}$, with $\epsilon>0$ small.
@@
By contradiction, there exists a $n_{0}$ tangency on the line $a_{n}-E{\lambda^{\theta n}}$ with $E=C-\frac{\epsilon}{\mu^{n_{0}/2}}$, $\epsilon>0$ small. This means that $z_{3,y}\mu^{n_{0}}=2$. By (4.6) and Lemma 7,
$z_{2,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\mu^{n}(a-a_{n}(t))+O\left(\mu^{n}\lambda^{2\theta n}\right)$
and by (4.19)
$z_{3,y}=Q\left(z_{2,y}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right),$
Hence,
$z_{3,y}\mu^{n_{0}}$ $=$ $Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$ (5.7)
$=$ $Q\frac{\epsilon^{2}}{\mu_{0}^{n_{0}}}\left(\lambda^{\theta}\mu\right)^{2n}\mu_{0}^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
$=$ $Q{\epsilon^{2}}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$
Since $\lambda^{\theta}\mu<1$, for $\epsilon$ small enough, this last quantity cannot be equal to $2$. The proof is completed. ∎
The following lemma shows that one cannot find a $n_{0}$ tangency on the line $a_{n}+\frac{T}{\mu^{n+n_{0}/2}}$ with $T$ a large constant.
Lemma 16.
For $T$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}+\frac{T}{\mu^{n+n_{0}/2}}$, with $T$ large.
@@
As in the proof of the previous lemma, by (4.6) and Lemma 7,
$z_{2,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\mu^{n}(a-a_{n}(t))+O\left(\mu^{n}\lambda^{2\theta n}\right)$
and by (4.19)
$z_{3,y}=Q\left(z_{2,y}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right)$ $=$ $Q\left(C\left(\lambda^{\theta}\mu\right)^{n}+\frac{T}{\mu^{n_{0}/2}}\right)^{2}+O\left(\frac{1}{\mu^{n}}\right)$
$>$ $Q\frac{T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n}}\right).$
Hence,
$z_{3,y}\mu^{n_{0}}$ $>$ $QT^{2}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$ (5.8)
For $T$ large enough, this last quantity cannot be equal to $2$, so in particular one cannot have a $n_{0}$ tangency. The proof is completed. ∎
5.3 Global curves of secondary tangencies
In the next proposition we prove the existence of global curves of secondary tangencies contained in $\left([-t_{0},t_{0}]\times\mathbb{R}\right)\cap\mathcal{B}_{n}$, see Figure 6.
Proposition 4.

### v3

@@ 5.1 Persistence of secondary tangencies
$D\Psi({0,0,0})=\left(\begin{matrix}0&\left(\lambda^{\theta}\mu\right)^{n}\Psi_{1,2}&\left(\lambda^{\theta}\mu^{2}\right)^{n}\Psi_{1,3}\\ \left(\frac{\mu^{3}}{\lambda^{2-2\theta}}\right)^{n}\Psi_{2,1}&\Psi_{2,2}&\Psi_{2,3}\end{matrix}\right).$ (5.1.2)
In particular, by (4.2), (2.2.3), (4.2), Lemma 2.2.5, (4.1.49) (in the estimation of the order term)
$\Psi_{1,2}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial t}-\frac{\partial W_{n-n_{0}}}{\partial t}\right]$ (5.1.3)
$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta}\mu\right)^{n}\right\}$
$+$ $O\left(n\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\right)-\frac{1}{\left(\lambda^{\theta}\mu\right)^{n}}\frac{\partial W_{n-n_{0}}}{\partial t}$
$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta}\mu\right)^{n}\right\},$
and by (4.2.31), (2.2.3), (4.2), Lemma 2.2.5 and (4.1.49) (in the estimation of the order term)
$\Psi_{1,3}$ $=$ $\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left[\frac{\partial m_{6,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial x}\frac{\partial m_{6,x}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}\right]$ (5.1.4)
$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left\{\mu^{n}+\frac{1}{DS_{C}}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}$
$+$ $O\left(\frac{\mu^{n}}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)+O\left(\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\frac{n-n_{0}}{\mu^{n-n_{0}}}\right)$
$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left[1+O\left(\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\right)\right]+O\left(\left(\lambda^{1-\alpha+\theta}\mu\right)^{n}\right)$
$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left[1+O\left(\frac{1}{\left(\lambda^{\theta}\mu^{2}\right)^{n}}\right)\right]+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)$
$=$ $\frac{DS_{C}}{\left(\lambda^{\theta}\mu\right)^{n}}\left[1+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right],$
@@
where we used that $\alpha<1$ and $\lambda^{1-\alpha+\theta}\mu<\lambda^{\theta}\mu$. Hence, by the expression for $S_{C}$, see Proposition 4.2.2, $\Psi_{1,2},\Psi_{1,3}>0$. Moreover, by Lemma 4.2.36, (4.2) and (2.2.4) we have $\Psi_{2,1}\neq 0$. This implies that $D\Psi({0,0,0})$ is onto and we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$. Moreover the tangent space of $\Psi$ in zero satisfies $T_{(0,0,0)}\Psi^{-1}(0)=\text{Ker}D\Psi$. Hence,
@@
$\frac{db}{dt}=-\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}=-\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}-{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right).$ (5.1.5)
In particular $db/dt$ is a negative number and $b$ is the graph of a smooth function. Furthermore, by the previous estimate on the slope of $b$, and by (2.2.7), the curve $b$ have a angle satisfying the following,
$\frac{db}{dt}=\frac{da_{n}}{dt}-{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right),$ (5.1.6)
where $\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]$ is away from zero, see (2.3.3).
Let $(t_{n_{0}},a_{n_{0}})$ be any parameter in $\mathcal{B}_{n}$ corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$. The proof is exactly the same as the previous case. In particular, by (4.2), (2.2.3), (4.2), (4.1.49) (in the estimation of the order term)
$\Psi_{1,2}$ $\geq$ $D(C-E)\frac{n}{\mu}\frac{\partial\mu}{\partial t}\left[1+O\left(\frac{1}{n\mu^{\theta n/2}}\right)\right],$
and by (4.2.25), (2.2.3), (4.2), Lemma 2.2.5 and (4.1.49) (in the estimation of the order term)
$\Psi_{1,3}$ $\geq$ $D(C-E)\left[1+O\left(\frac{1}{\mu^{\theta n/2}}\right)\right].$
Moreover, by (4.2.27), (2.2.3), (4.2),
$\Psi_{1,2}$ $\leq$ $\frac{DT}{\left(\lambda^{\theta}\mu\right)^{n}\mu^{n_{0}^{*}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}\left[1+O\left(\frac{1}{\mu^{n_{0}^{*}/2}}\right)\right].$
Since, by Proposition 4.2.2 and Remark 2.1.6, $D,C-E,T,{\partial\mu}/{\partial t}>0$, then $\Psi_{1,2}>0$, $\Psi_{1,3}>0$ and $\Psi_{2,1}>0$, see Lemma 4.2.36. As before, because $D\Psi({0,0,0})$ is onto we get that the set of secondary tangencies $\Psi^{-1}(0)$ is locally the graph of a ${\mathcal{C}}^{2}$} function $b$. Moreover,
$\left|\frac{db}{dt}\right|=\left|\frac{\Psi_{1,2}}{\Psi_{1,3}}\frac{1}{\mu^{n}}\right|=O\left(\frac{n}{\left(\lambda^{\theta}\mu^{2}\right)^{n}\mu^{n_{0}^{*}/2}}\right),$ (5.1.7)
which goes to zero exponentially fast, see (2.3.1). The proof is complete. ∎
5.2 Properties of the loci of secondary tangencies
We prove now that for any $n_{0}$ by moving in the direction of the $a$ parameter one can find at the most one tangency of type $n_{0}$.
Lemma 5.2.1.
@@ 5.2 Properties of the loci of secondary tangencies
$\frac{\partial z_{3,y}}{\partial a}-\frac{\partial W_{n-n_{0}}}{\partial a}>0.$
@@
Using the estimates in Proposition 4.2.2, one notices that the leading term of $\partial z_{3,y}/{\partial a}$ is proportional to $\left(\lambda^{\theta}\mu^{2}\right)^{n}$ which is strictly larger than one, while $\partial W_{n-n_{0}}/{\partial a}$ is strictly smaller than one, see Lemma 2.2.5, for $n$ large enough. The inequality follows. ∎
@@
Lemma 5.2.2.
For all $\epsilon>0$ and for $n$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}-E{\lambda^{\theta n}}$ where $E=C-\frac{\epsilon}{\mu^{n_{0}^{*}/2}}$.
@@
By contradiction, there exists a $n_{0}$ tangency on the line $a_{n}-E{\lambda^{\theta n}}$ with $E=C-\frac{\epsilon}{\mu^{n_{0}^{*}/2}}$. This means that $z_{y}^{(3)}\mu^{n_{0}}=2+O(\lambda^{n_{0}})$. By Lemma 4.1.23,
$z_{y}^{(3)}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+z_{y}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+a.$
where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)=(C-E)\left(\lambda^{\theta}\mu\right)^{n}\left(1+O\left({n}/{\mu^{n}}\right)\right)$. Hence, by (2.3.1), (4.0.1) and Lemma 4.1.48
$z_{y}^{(3)}\mu^{n_{0}}$ $=$ $Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$ (5.2.3)
$=$ $Q\frac{\epsilon^{2}}{\mu^{n_{0}^{*}}}\left(\lambda^{\theta}\mu\right)^{2n}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
$=$ $O\left(\left(\lambda^{\theta}\mu\right)^{2n}\mu^{\alpha n-\frac{1}{2}\alpha n}\right)+O\left(\frac{1}{\mu^{n-n_{0}}}\right)$
$=$ $O\left(\left(\lambda^{3\theta}\mu^{7/4}\right)^{2n}\right).$
Since $\lambda^{3\theta}\mu^{7/4}<1$, see (2.3.1), this last quantity cannot be close to $2$. Contraddiction.The proof is complete. ∎
The following lemma shows that one cannot find a $n_{0}$ tangency on the line $a_{n}+\frac{T}{\mu^{n+n_{0}^{*}/2}}$ with $T$ a large constant.
Lemma 5.2.4.
For $T$ large enough, there are no secondary tangencies of type $n_{0}$ along the line $a_{n}+\frac{T}{\mu^{n+n_{0}^{*}/2}}$.
@@
As in the proof of the previous lemma, By Lemma 4.1.23,
$z_{y}^{(3)}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+z_{y}(t,a)=QH^{2}\left[1+O\left(H\right)\right]+a.$
where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)=\frac{T}{\mu^{n_{0}^{*}/2}}\left(1+O\left(\left(\lambda^{\theta}\mu\right)^{n}\mu^{n_{0}^{*}/2}\right)\right)$. Hence, by (2.3.1), and Lemma 4.1.48
$z_{y}^{(3)}\mu^{n_{0}}>\frac{QT^{2}}{2\mu^{n_{0}}}\mu^{n_{0}}+O\left(\frac{1}{\mu^{n}}\mu^{n_{0}}\right)=\frac{QT^{2}}{2}+O\left(\frac{1}{\mu^{n-n_{0}}}\right).$ (5.2.5)
For $T$ large enough, this last quantity cannot be close to $2$, so in particular one cannot have a $n_{0}$ tangency. The proof is complete. ∎
5.3 Curves of secondary tangencies with uniform size
In the next proposition we prove the existence of curves of secondary tangencies which are graphs of functions defined over the full domain $t_{0},t_{0}]$, see Figure 6.
Proposition 5.3.1.